wiki:UC20170410

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 4/10/17

Attending: Nancy Chanover, Rachel Kuzio de Naray, Scott Anderson, Bill Ketzeback, Mark Klaene, John Tobin (for John Wisniewski), Mary Beth Kaiser, Joanne Hughes, Jason Lozo, Jon Holtzman, Rene Walterbos, John Bally, Ben Williams, Sarah Tuttle, Anne Verbiscer, Chip Kolbulnicky, and Bruce Gillespie


User feedback and comments from institutional representatives:

  • Washington - Scott had nothing new to report.
  • Georgia State - Rachel had nothing new to report.
  • Seattle - Joanne had nothing new to report.
  • NMSU - Jon had nothing new to report.
  • JHU - Mary Beth said that the 3.5-m wiki page seems to need a non-VPN user connection. Jon Holtzman said that you don’t need to connect to port 8000 because it gets redirected — this might help those logging on from VPN environments.
  • Oklahoma - John Tobin said that students had a great class trip to APO, except for the weather.
  • Colorado - John Bally had nothing new to report.
  • Wyoming - Chip had nothing new to report.
  • Virginia - Anne had nothing new to report.
  • NAPG - No report.
  • Adler - No report.

Telescope and Instruments Report:

Mark noted that we’ve suffered through some late spring dust and snow storms. Otherwise, the observing systems have been nearly trouble-free during the past month. Mark noted that Jason Lozo is leaving in order to make a career move — we’ll miss him and are grateful for his contributions. Virginia is sending a host of experts to grapple with the TripleSpec readout anomalies, starting on 15 May. For ARCSAT, major upgrades were made to the control software, and we are now cleaning up a few loose ends. Lastly, a new site-wide phone system was installed and brought on line. Mark’s detailed report follows below:

3.5-m Telescope and Instruments Highlights, 3/9/17 through 4/5/17

0) Overview

Observing time has been interrupted with spring winds, clouds, dust, and a late winter storm that brought some nice snow to the mountain but overall fairly productive period. Jason Lozo, the engineer at UW that has been working on some 3.5-m and ARCSAT issues, will be leaving to take a new job. The impact of this to the 3.5-m is still being discussed.

1) Telescope - Telescope was operational with only minimal issues

2) Instruments

DIS blue and red cameras continue to be operational along with slitviewer. The red camera shows some indication of return of contamination but still at a minimal amount. We continue to monitor frequently and have temperature humidity readings inside instrument as well.

TripleSpec is warm and ready for a servicing trip by UVa instrument personnel starting May 15th.

Agile has been operational this period.

Echelle has been operational with no problems. We are looking at April 19th through May 1st for vacuum servicing.

NIC-FPS is operational.

ARCTIC is operational. The diffuser is available without spinning. A replacement for the damaged diffuser is in progress. Jason is set up to finalize his pneumatic shutter on the bench.

3) ARCSAT - Most, but not yet quite all, of the functions on ARCSAT are restored with a new software upgrade. Some issues will continue as expected since we are using commercial software.

4) Site News

APO is completing the final phase of a site phone system upgrade. Please see the following link if your users plan to come to the site: http://www.apo.nmsu.edu/Site/Computing_Resources/Phone-System/APO_Phone_System.html


Q2 3.5m scheduling

Ben said that the ARCSAT schedule has been posted; some Q2 half-nights are still open. Mark mentioned that his staff will support ARCSAT instrument changes for Memorial Day by changing the instruments either on the Friday before or the Tuesday after. There was no news concerning the 3.5m Q2 schedule.


Report on DIS upgrades PDR

Nancy and Sarah are nearly finished writing up the report on the DIS upgrades PDR. There is a clear plan for work to be performed in September (which is not concurrent with the summer shutdown for the telescope). Feedback is still welcome from users regarding the impact to their science programs resulting from the red camera contamination.


Status of the schedule for summer shut down

Mark said there are no changes to the plans for the summer shutdown (so far). The shutdown dates are start 31 July and return to science on 18 August, weather permitting. The primary will get washed, and M2 and M3 will be realuminized.


Fiber feed project from 3.5-m to APOGEE

Nancy said that with several instrumentation and infrastructure projects underway, she would like to revisit the priorities of the user community so that the limited engineering resources can be appropriately applied. As previously mentioned, the fiber feed project was described in a white paper (see Instruments); the paper is now about two years old. Nancy wants to reopen the discussion to ascertain the level of user interest and scientific potentials in implementing the fiber feed system. This is important now because we have new ARC partners and several new telescope users who were not present when this project was launched.

ACTION ITEM: Users Committee members to reassess level of user interest and scientific potentials for fiber feed system between the 3.5-m telescope and the APOGEE spectrograph.

Mary Beth said that one of the issues in the system design was relatively high costs. Jon said that we need to assess user demand and potential science capabilities again, independent of the cost question. Mark is in favor of the simpler implementation of the fiber pickoff at the NA2 port, rather than outfitting an unused port with a new guider and rotator. After an extended discussion, Nancy said that she wants to know if there is a reasonable number of committed potential users at most of the ARC institutions. See the white paper at Instruments for a description of the capability options.


Miscellaneous

Mark announced that Dan Long will be retiring on 28 April, following a most excellent and legendary 25-year career at APO.


ACTION ITEMS:

open from previous meetings: None

new from this meeting:

  • Users Committee members to reassess level of user interest and scientific potentials for fiber feed system between the 3.5-m telescope and the APOGEE spectrograph. See Instruments for detailed information on this project.

Next meeting:

The next Users Committee phonecon will be on Monday 08 May at 1030 MDT — the agenda and other materials will be sent to committee members during the preceding week.

Note: See TracWiki for help on using the wiki.